# Safe-CLIP: Removing NSFW Concepts from Vision-and-Language Models

### v2

@@ Appendix 0.D Discussion and Limitations
Our fine-tuning strategy would not be effective if the model did not acquire knowledge of inappropriate concepts during pre-training. Therefore, we do not advise removing inappropriate content entirely from the training data but rather conceive our approach as a more general post-training strategy that could be applied before model release to remove the impact of inappropriate concepts.
Our fine-tuning strategy is based on the collection of toxic content, predicted from an LLM fine-tuned to generate inappropriate content. We realize that this model has strong and direct ethical implications, as such we commit not to release the model by any means. Further, our methodology might have further ethical implications. For instance, one could construct a model generating exclusively inappropriate content by reversing the fine-tuning losses. However, it could also be applied to cases beyond inappropriateness, and as a general framework for cleaning generative and discriminative models. Finally, the model representation of inappropriateness and toxic content can reflect the societal dispositions of the social groups represented in the training data of Llama 2 and in the ViSU dataset and might lack a more diverse sentiment.
Limitations. While our model can remove the impact of inappropriate concepts in a wide variety of cases, it does not provide any guarantee of success. For instance, it might fail to remove inappropriate content under certain conditions. Some failure cases are reported in Fig. 14. Future works might want to further reduce the impact of these failure cases by enlarging the training dataset in both quantity and variance. Warning: The following pages contain blurred texts and images due to inappropriate content.
@@
Figure 13: Additional sample captions generated from NSFW images with LLaVA and LLaVA + Safe-CLIP.

### v3

@@ Appendix 0.D Discussion and Limitations
Our fine-tuning strategy would not be effective if the model did not acquire knowledge of inappropriate concepts during pre-training. Therefore, we do not advise removing unsafe content entirely from the training data; rather, we propose our approach as a more general post-training strategy that could be applied before the model is released to remove the impact of inappropriate concepts.
Our fine-tuning strategy is based on the collection of toxic content, predicted from an LLM fine-tuned to generate inappropriate content. We realize that this model has strong and direct ethical implications, as such we commit not to release the model by any means. Further, our methodology might have additional ethical implications, as the model’s representation of inappropriateness and toxic content can reflect the societal dispositions of the social groups represented in the training data of Llama 2 and in the ViSU dataset. This, in turn, might result in a lack of more diverse sentiments.
Addressing the legality of the dataset. The dataset employed in our research adheres to all pertinent legal standards and ethical guidelines. Specifically, the safe images come from the publicly available COCO dataset, which is well-established in the literature and legally compliant for research purposes. Regarding the NSFW images, these are synthetically generated using a publicly available diffusion model from Hugging Face, ensuring that no real individuals are depicted, thus eliminating privacy concerns. The NSFW images fall into the seven categories of inappropriate content previously defined in the literature [44] (i.e. hate, harassment, violence, self-harm, sexual, shocking, and illegal activities). These categories ensure comprehensive coverage of potentially inappropriate content that our model aims to filter out.
The textual data follows a similar ethical protocol. All real text data is derived from the COCO dataset, ensuring that it is ethically sourced and legally compliant. The NSFW textual data are generated using a fine-tuned version of the Llama 2 model, starting from the safe textual sentences contained in the COCO dataset. This approach involves modifying the safe sentences to introduce NSFW elements deliberately. The intent is to create controlled instances of harmful or unethical content, which are essential for training and fine-tuning our model to recognize and filter out inappropriate content effectively. This method allows us to build a robust system capable of maintaining ethical and legal compliance in real-world applications.
As reported in the main paper, the ViSU dataset is released in a controlled manner, with only the textual portion available to verified researchers. Access requires a declaration of research-only use, preventing malicious purposes. Due to its explicit nature, NSFW images are not publicly released to avoid potential misuse, while still ensuring their reproducibility for validation by the research community. The dataset containing NSFW images is securely stored, with access restricted solely to the researchers of this project, ensuring strict control over sensitive content.
We believe that by carefully generating and utilizing NSFW content within a controlled and ethical framework, we contribute to the development of safer and more responsible AI systems. Our work is guided by a commitment to ethical research practices and compliance with all legal standards, ensuring that our contributions align with broader societal and ethical goals.
Limitations. While our model can remove the impact of inappropriate concepts in a wide variety of cases, it does not provide any guarantee of success. For instance, it might fail to remove inappropriate content under certain conditions. Some failure cases are reported in Fig. 13. Future works might want to further reduce the impact of these failure cases by enlarging the training dataset in both quantity and variance. Warning:
The following pages contain blurred texts and images due to inappropriate content.
@@
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Figure 12: Sample captions generated from NSFW images with LLaVA and LLaVA + Safe-CLIP.